1. Calculate the pH of a buffer containing 1.0 M CH3COOH and 1.0 M CH3COONa. (Ka = 1.8 ✕ 10–5)
1
Expert's answer
2020-11-09T14:04:33-0500

PH = Pka - log (acid/salt)

PH = -log(1.8×105) - log (1/1)

PH= 4.74-0

PH= 4.74